A bucket contains 10 balls labeled as follows: A,A,B,C,C,C,C,D,E,E
What is the probability of drawing a ball labeled C? Express the probability as a decimal.
What is the probability of drawing a ball labeled with something other than a B? Express the probability as a percent.
What is the probability of drawing an A,replacing it and then drawing A? Express the probability as a fraction in simplest form.
What is the probability of drawing an E, NOT replacing it and then C. Express the probability as a percent
What is the probability of drawing an A or B? Express the probability as a decimal.

Answer by jim_thompson5910(35256)

A bucket contains 10 balls labeled as follows: A,A,B,C,C,C,C,D,E,E
What is the probability of drawing a ball labeled C? Express the probability as a decimal.
What is the probability of drawing a ball labeled with something other than a B? Express the probability as a percent.
What is the probability of drawing an A,replacing it and then drawing A? Express the probability as a fraction in simplest form.
What is the probability of drawing an E, NOT replacing it and then C. Express the probability as a percent
What is the probability of drawing an A or B? Express the probability as a decimal.
------------------------------------------------------------------------------------------------------------------------------
------------------------------------------------------------------------------------------------------------------------------



What is the probability of drawing a ball labeled C? Express the probability as a decimal.

There are 4 C's out of 10 total.

4/10 = 0.4
-------------------------------------------------------

What is the probability of drawing a ball labeled with something other than a B? Express the probability as a percent.


There's only 1 B out of 10 total. So 10-1 = 9 balls are not B.

9/10 = 0.9 = 90%
-------------------------------------------------------

What is the probability of drawing an A,replacing it and then drawing A? Express the probability as a fraction in simplest form.


P(A) = 2/10 = 1/5

P(A and A) = P(A)*P(A)
P(A and A) = (1/5)*(1/5)
P(A and A) = 1/25
-------------------------------------------------------

What is the probability of drawing an E, NOT replacing it and then C. Express the probability as a percent


P(E then C) = (2/10)*(4/9) = 8/90 = 4/45 = 0.08889 = 8.889 %
-------------------------------------------------------

What is the probability of drawing an A or B? Express the probability as a decimal.


P(A or B) = (# of balls that are either A or B)/(# total)
P(A or B) = 3/10
P(A or B) = 0.3